# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ IV.2 The geometry of ${\sf BS}$ and the fragmentation of ${\sf Dyn}_{\sf BS}$
where the numerical results in Fig. 4—obtained by sampling random words and post-selecting on membership in $K_{e,L}$—indicate that $\alpha\approx 1.84$. Models with ${\sf Dyn}_{\sf BS}$ dynamics provide an example (indeed the first that we are aware of) of a strongly fragmented model where the largest sector occupies neither an exponentially nor polynomially small fraction of Hilbert space.
@@ IV.3 The Dehn function
$w_{\mathrm{large}}(n)\triangleq\mathtt{a}^{-1}\mathtt{b}^{-n}\mathtt{a}^{-1}\mathtt{b}^{n}\mathtt{a}\mathtt{b}^{-n}\mathtt{a}\mathtt{b}^{n}$ (70)
which is of length $|w_{\mathrm{large}}(n)|=4n+4$, belongs to $K_{e,L}$, and is shown in Fig. 3 as the thick dashed line in panel (b). As a path, this word can be broken into two legs. The first leg moves to the vertex labelled by $\mathtt{a}^{2^{n}}$ by first going “up” the hyperbolic plane of a given sheet, traversing one step along ${\mathbf{\hat{x}}}$, and then coming back “down”. The second leg moves back to the origin, but does so by passing along a different sheet of the tree. From Fig. 3, it is clear that the area of the minimal surface bounding $w_{\mathrm{large}}(n)$ is exponentially large; more precisely it is
@@
which scales exponentially with $n$. This construction thus shows that ${\sf Dehn}(L)=\Omega(2^{L})$ Gersten (1992). In App. E we show that in fact ${\sf Dehn}(L)=\Theta(2^{L})$. From the results of Sec. III, this implies that ${\sf Dyn}_{G}$ has exponentially long mixing, relaxation, and hitting times. A more sophisticated treatment needs to be given in order to understand the scaling of the typical Dehn function $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$. To our knowledge, this question has not been answered in the math literature. While we will not provide a completely rigorous proof of $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$’s scaling, a combination of physical arguments and numerics—which we relegate in their entirety to App. E—indicate that
@@ IV.4 Exponentially slow hydrodynamics
with probability 1 in the large $n$ limit. The long relaxation time of the density wave is attributed to the long mixing time of ${\sf BS}$ because of its large Dehn function (see Sec. III). This is remarkable because probing the large scale complexity of ${\sf BS}$ only requires studying the dynamics of local operators, viz. those that overlap with $n_{\mathtt{b}}$. We henceforth will denote the relaxation and mixing times by $t_{th}$. Of course since $2^{n}$ is the smallest possible time needed to map $w_{\rm large}(n)$ to $\mathtt{e}^{4n+4}$, it gives only a lower bound on $t_{\rm th}$. In the case where ${\sf Dyn}_{\sf BS}(t)$ is generated by a classical Markov process or random unitary circuit dynamics, ${\sf Dyn}_{G}(t)$ effectively leads to words executing a random walk on configuration space of loops. Due to the diffusive nature of this random walk, this leads us to expect that the true thermalization time instead scales as the square of its lower bound, viz.
@@
and that $|w_{A,q}\rangle$ is chosen randomly from the set of all states in $K_{e,L}$ that have this expectation value (with the restriction to $K_{e,L}$ done purely for notational simplicity). The “depth” $n_{A}$ that $w_{A,q}$ proceeds into ${\mathsf{CG}}_{\sf BS}$ is equal to the contrast of the density wave, which we define as
[…]
@@
Note that this exponential timescale is not visible in the standard linear-response limit, in which one takes $A\to 0$ before taking $q\to 0$: the two limits lead to qualitatively different relaxation. In the standard linear-response limit, fluctuations at scale $q$ will decorrelate on the typical relaxation timescale $\sim 2^{\sqrt{L}}$. Beyond $t_{\rm th}$, we would also like to know how the amplitude $A$—or equivalently the contrast $n_{A}$—behaves as a function of time. To estimate this, consider first how $w_{\rm large}(n)$ relaxes. By the geometric considerations of Sec. IV.2, the shortest homotopy reducing $w_{\rm large}(n)$ to the identity word must first map $w_{\rm large}(n)\rightarrow w_{\rm large}(n-1)$, then $w_{\rm large}(n-1)\rightarrow w_{\rm large}(n-2)$, and so on; see Fig. 8 for an illustration. The self-similarity of this process means that the time needed to map $w_{\rm large}(m)\rightarrow w_{\rm large}(m-1)$, viz. $t_{\rm th}(w_{\rm large}(m-1))$, is equal to the combined time needed to perform all of the maps $w_{\rm large}(k)\rightarrow w_{\rm large}(k-1)$ with $k<m$. Applying this observation to the density wave under consideration and using (74), we estimate
@@
as the time at which the collapse of the density wave becomes noticeable within a precision controlled by $10<\varepsilon<$ 1. Then the form (79) implies that
@@
implying that the collapse is instantaneous in the large-contrast limit. This behavior is strongly reminiscent of the cutoff phenomenon observed in many types of Markov chains (see e.g. Aldous and Diaconis (1986); Diaconis (1996)), and it could be interesting to explore this connection further in future work.
@@ IV.5.2 Slow relaxation
Fig. 5 (c) shows the scaling of $t_{\rm th}$ with $L$, which is observed to admit an excellent fit to the predicted scaling of $\sim 2^{2n}$. While this confirms that the density wave present in $|w_{\rm large}(n,L)\rangle$ relaxes on a timescale of $t_{\rm th}\sim 2^{2n}$, it does not show that all states with $\mathtt{b}$ density waves of amplitude $A$ and wavelength $q$ relax on times of order $2^{2A/q}$. This in fact cannot be true, and fast-relaxing density waves always exist. Indeed, consider the word $w_{\mathrm{small}}(n,L)$ obtained from $w_{\mathrm{large}}(n,L)$ by replacing all $\mathtt{a}$s and $\mathtt{a}^{-1}$s with $\mathtt{e}$s. The Dehn time of this word is merely ${\sf Dehn}(w_{\mathrm{small}}(n,L))\sim n^{2}$, since there are no $\mathtt{a}$s present to slow down the dynamics of the $\mathtt{b}$s (any $\mathtt{a}\mathtt{a}^{-1}$ pairs created in between the segments of the density waves have a net zero number of $\mathtt{a}$s, and are thus ineffectual at providing a slowdown). This means that the relaxation time of an initial state $|w\rangle$ carrying a $n_{\mathtt{b}}$ density wave cannot be predicted from knowledge of the conserved density alone—one must also have some knowledge about the distribution of $\mathtt{a}$’s in $|w\rangle$. We expect however that $t_{\rm th}\sim 2^{2A/q}$ for generic states containing an amplitude-$A$, wavelength-$q$ density wave. Indeed, as argued above near (73), this is simply because as long as the value of the $\mathtt{a}$-charge $\sum_{i}(|\mathtt{a}\rangle\langle\mathtt{a}|_{i}-|\mathtt{a}^{-1}\rangle\langle\mathtt{a}^{-1}|_{i})$ is nonzero in the regions between the segments of the density wave—which will almost always be true for a random density wave state in the thermodynamic limit—the $\mathtt{a}$s trapped “inside” the density wave will take an exponentially long time to escape. Note that this remark applies to a generic density wave state $|w\rangle$, regardless of whether or not $|w\rangle\in K_{e,L}$. In Fig. 6, we numerically investigate the relaxation of generic density waves by considering initial states which host density waves of momentum $q=4\pi/3L$ and amplitude $A=nq$ (with $n$ fixed at $n=L/10$), but which are otherwise random. Compared with $|w_{\rm large}(n,L)\rangle$, these random density waves (which generically lie in different $K_{g,L}$ sectors) exhibit a much broader range of thermalization times; some thermalize very quickly, while some never thermalize over our chosen simulation time window (Fig. 6 a). Nevertheless, we still observe the average thermalization timescale to scale exponentially with $n$ (Fig. 6 b). The large sample-to-sample fluctuations of $t_{\rm th}$ make it difficult to reliably extract the exact scaling behavior, but the above reasoning suggests that $t_{\rm th}$ continues to scale as $2^{2n}$ for typical initial states.
@@
Figure 6: Relaxation of $n_{\mathtt{b}}$ under stochastic circuit dynamics for initial states with random $\mathtt{b}$ density waves. The initial states are chosen to be words of the form $w_{1}\mathtt{b}^{n}w_{2}\mathtt{b}^{-n}w_{3}\mathtt{b}^{n}w_{4}\mathtt{b}^{-n}w_{5}$, where $n=L/10$ and the $w_{1,\dots,5}$ are random words containing only the characters $\{\mathtt{a},\mathtt{a}^{-1},\mathtt{e}\}$. The time averaging window is $T=100$. (a) Time evolution of observable $\overline{\expectationvalue{n_{\mathtt{b}}}^{2}}$ from Eq. (85) for several independent runs with random initial states described above. The time at which the density wave collapses and the final equilibrium value of $\mkern 1.5mu\overline{\mkern-1.5mu\langle n_{\mathtt{b}}\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$ are seen to change significantly for different choices of initial state, with some runs thermalizing faster than $T$ (low-lying blue lines) and some runs not thermalizing within the whole displayed time window of $5000T$ (high-lying blue lines). (b) The thermalization time $t_{\rm th}$ of the density wave for initial states that exhibit a significant drop in $\mkern 1.5mu\overline{\mkern-1.5mu\langle n_{\mathtt{b}}\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$ during the displayed time window. $t_{\rm th}$ is seen to scale exponentially with $n$, although the significant sample-to-sample fluctuations prevent a reliable extraction of the base of the exponent (the best-fit line in the plot is drawn to $t_{\rm th}\sim 2^{1.37n}$).
@@ V Fragile fragmentation and the space complexity of the word problem
Our discussion in the past few sections has focused on the way in which the time complexity of the word problem enters in the thermalization times of ${\sf Dyn}_{G}$. In this section, we turn to the space complexity of the word problem. As discussed in Sec. II.2, the space complexity of the word problem is determined by the maximal amount of space required to map between two words $w,w^{\prime}\in K_{g,L}$, as diagnosed by the expansion length function ${\sf EL}_{g}(L)$ (22). When the expansion length is large, transitioning from $w$ to $w^{\prime}$ necessarily requires that $w$ first grow to be much larger than its original size before shrinking down to $w^{\prime}$. When ${\sf EL}_{g}(L)>L$, the dynamics thus lacks the spatial resources needed to connect all states which describe the same group element. In this situation, ${\sf Dyn}_{G}$ cannot act ergodically in $K_{g,L}$, and thus the $K_{g,L}$ become further fragmented. Each fragment now contains words that can be reached from some reference word $w$ by derivations that do not involve intermediate words longer than $L$. We call this phenomenon fragile fragmentation in analogy with the notion of fragile topology in band theory Po et al. (2018): ${\sf Dyn}_{G}$ is said to exhibit fragile fragmentation if there are pairs of words $w,w^{\prime}$ of length $L$ such that ${\sf Dyn}_{G}$ on a system of length $L$ does not connect |$\ket{w}$ and |$\ket{w^{\prime}}$, but ${\sf Dyn}_{G}$ on a larger system of length $L^{\prime}>L$ connects the “padded” words |$\ket{w}\otimes\ket{\texttt{e}^{L^{\prime}-L}}$ and |$\ket{w^{\prime}}\otimes\ket{\texttt{e}^{L^{\prime}-L}}$. [^17] For semigroups, we would replace e with a character 00 which freely moves past other characters. A schematic illustration of this definition is given in Fig. 7. A simple example of this phenomenon that exists in higher dimensions is the jamming transition. In jammed systems, an ensemble of particles with hard core repulsive interactions can exhibit a phase transition from a low density mobile phase to a high density jammed phase. The analog of fragmentation is the limited configuration space that particles can explore in the jammed phase. When the jammed particles are given more space, their density decreases, and when it drops below a critical value, the dynamics becomes ergodic. If the extra space is subsequently removed ergodicity may again be broken, but the system may find itself in a previously inaccessible microstate. The models we discuss in this section are more drastic examples of this phenomenon: unlike the examples with hardcore particle models, the models we study exihibit jamming even in one dimension, where the analog of the critical jamming density can be polynomially or exponentially small in system size. To understand when ${\sf Dyn}_{G}$ exhibits fragile fragmentation, we need to compute the expansion length ${\sf EL}(L)$. [^18] Our notation in this discussion will assume that $G$ is a group, for which ${\sf EL}(L)\triangleq{\sf EL}_{e}(L)\gtrsim{\sf EL}_{g}(L)$ for all $g$. For general semigroup dynamics, one would need to consider the ${\sf EL}_{g}(L)$’s separately. Doing so for a general group can be rather difficult, as ${\sf EL}(L)$’s definition involves a rather complicated minimization problem. If however we already know the time complexity of the word problem—i.e. if we know the scaling of ${\sf Dehn}(L)$—it is possible to place a lower bound on ${\sf EL}(L)$ Gersten and Riley (2002). Indeed, suppose a word $w\in K_{e,L}$ has an expansion length ${\sf EL}(w)$, implying that $|w|$ is at most ${\sf EL}(w)$ during any homotopy from $w$ to $\mathtt{e}^{L}$. Then the expansion length of any derivation $D(w\rightsquigarrow\mathtt{e})$ cannot exceed the total number of length-${\sf EL}(w)$ words; if it did, there would be at least one state which appeared multiple times in $D(w\rightsquigarrow\mathtt{e})$—which implies that such a derivation cannot be of minimal length. Since the number of length ${\sf EL}(w)$ words is $(2|S|+1)^{{\sf EL}(w)}$, we thus have ${\sf Dehn}(w)\leq(2|S|+1)^{{\sf EL}(w)}$. By maximizing over all possible $w\in K_{e,L}$ and taking a logarithm, we obtain the general bound

### 11 Apr 2024 v2

@@ IV.2 The geometry of ${\sf BS}$ and the fragmentation of ${\sf Dyn}_{\sf BS}$
where the numerical results in Fig. 4—obtained by sampling random words and postselecting on membership in $K_{e,L}$—indicate that $\alpha\approx 1.84$. Models with ${\sf Dyn}_{\sf BS}$ dynamics provide an example (indeed the first that we are aware of) of a strongly fragmented model where the largest sector occupies neither an exponentially nor polynomially small fraction of Hilbert space.
@@ IV.3 The Dehn function
$w_{\mathrm{large}}(n)\triangleq\mathtt{a}\mathtt{b}^{-n}\mathtt{a}^{-1}\mathtt{b}^{n}\mathtt{a}^{-1}\mathtt{b}^{-n}\mathtt{a}\mathtt{b}^{n}$ (73)
which is of length $|w_{\mathrm{large}}(n)|=4n+4$, belongs to $K_{e,L}$, and is shown for $n=3$ in Fig. 3 as the thick dashed line in panel (b). As a path, this word can be broken into two legs. The first leg moves to the vertex labelled by $\mathtt{a}^{2^{n}}$ by first going “up” the hyperbolic plane of a given sheet, traversing one step along ${\mathbf{\hat{x}}}$, and then coming back “down”. The second leg moves back to the origin, but does so by passing along a different sheet of the tree. From Fig. 3, it is clear that the area of the minimal surface bounding $w_{\mathrm{large}}(n)$ is exponentially large; more precisely it is
@@
which scales exponentially with $n$. This construction thus shows that ${\sf Dehn}(L)=\Omega(2^{L})$ [56]. This bound is in fact tight (see App. E), and so
${\sf Dehn}(L)=\Theta(2^{L}).$ (75)
From the results of Sec. III, this implies that ${\sf Dyn}_{G}$ has exponentially long mixing, relaxation, and hitting times. A more sophisticated treatment needs to be given in order to understand the scaling of the typical Dehn function $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$. To our knowledge, this question has not been answered in the math literature. While we will not provide a completely rigorous proof of $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$’s scaling, a combination of physical arguments and numerics—which we relegate in their entirety to App. E—indicate that
@@ IV.4 Exponentially slow hydrodynamics
with probability 1 in the large $n$ limit. The long relaxation time of the density wave is attributed to the long mixing time of ${\sf BS}$ because of its large Dehn function (see Sec. III). This is remarkable because probing the large scale complexity of ${\sf BS}$ only requires studying the dynamics of local operators, namely those that overlap with $n_{\mathtt{b}}$. We henceforth will denote the relaxation and mixing times by $t_{th}$. Of course since $2^{n}$ is the smallest possible time needed to map $w_{\rm large}(n)$ to $\mathtt{e}^{4n+4}$, it gives only a lower bound on $t_{\rm th}$. In the case where ${\sf Dyn}_{\sf BS}(t)$ is generated by a classical Markov process or random unitary circuit dynamics, ${\sf Dyn}_{G}(t)$ effectively leads to words executing a random walk on configuration space of loops. Due to the diffusive nature of this random walk, this leads us to expect that the true thermalization time instead scales as the square of its lower bound, namely
@@
and that $|w_{A,q}\rangle$ is chosen randomly from the set of all states in $K_{e,L}$ that have this expectation value (with the restriction to $K_{e,L}$ done purely for notational simplicity). The “depth” $n_{A}$ that $w_{A,q}$ proceeds into ${\mathsf{CG}}_{\sf BS}$ is equal to the contrast of the density wave, which we define as the integral of the $\mathtt{b}$ density over a quarter period of the density wave:
$n_{A}\triangleq\int_{0}^{\pi/2q}n_{\mathtt{b}}(x)=\frac{A}{q}.$ (80)
@@
Note that this exponential timescale is not visible in the standard linear-response limit, in which one takes $A\to 0$ before taking $q\to 0$: the two limits lead to qualitatively different relaxation. In the standard linear-response limit, fluctuations at scale $q$ will decorrelate on the typical relaxation timescale $\sim 2^{\sqrt{L}}$. Beyond $t_{\rm th}$, we would also like to know how the amplitude $A$—or equivalently the contrast $n_{A}$—behaves as a function of time. To estimate this, consider first how $w_{\rm large}(n)$ relaxes. By the geometric considerations of Sec. IV.2, the shortest homotopy reducing $w_{\rm large}(n)$ to the identity word must first map $w_{\rm large}(n)\rightarrow w_{\rm large}(n-1)$, then $w_{\rm large}(n-1)\rightarrow w_{\rm large}(n-2)$, and so on; see Fig. 9 for an illustration. The self-similarity of this process means that the time needed to map $w_{\rm large}(m)\rightarrow w_{\rm large}(m-1)$, namely $t_{\rm th}(w_{\rm large}(m-1))$, is equal to the combined time needed to perform all of the maps $w_{\rm large}(k)\rightarrow w_{\rm large}(k-1)$ with $k<m$. Applying this observation to the density wave under consideration and using (78), we estimate
@@
as the time at which the collapse of the density wave becomes noticeable within a precision controlled by the constant $10<\varepsilon<$ 1. Then the form (83) implies that
@@
implying that the collapse is instantaneous in the large-contrast limit. These arguments show that density waves collapse abruptly, but do not tell us about the statistical distribution of collapse times obtained when considering an ensemble of realizations of the dynamics (specific disorder realizations of random unitary circuits, specific choices of Hamiltonian, etc.). In systems which relax suddenly, the width $\sigma_{t_{\rm th}}$ of this distribution can be either parametrically smaller (in system size) than the mean collapse time $t_{\rm th}$, or can scale as $\Omega(t_{\rm th})$. The former case occurs in Markov chains displaying a “cutoff” (see e.g. [38, 39]), which can occur when the chain describes random motion on a highly-connected graph, or when it describes relaxation in a metastable system (where relaxation is caused by rare events that yield a Poissonian scaling $\sigma_{t_{\rm th}}\sim\sqrt{t_{\rm th}}=o(t_{\rm th})$). The latter case is typical in systems which relax diffusively, with an unbiased random walk on $\mathbb{Z}^{d}$ being a typical example. From the geometric considerations of this section, it is perhaps not clear which scenario should apply a priori. The numerics of the following section tentatively point to the latter scenario, with the distribution of thermalization times roughly obeying $\sigma_{t_{\rm th}}\sim t_{\rm th}$. Giving a more precise characterization of this scaling would be interesting to explore in future work.
@@ IV.5.2 Slow relaxation
Figure 6: Distribution of thermalization times $t_{\rm th}$ of the density wave defined by $|w_{\rm large}(n,L)\rangle$ with $n=L/10$, shown for different system sizes with 300 circuit realizations. A rather broad distribution is observed.
We now investigate how the thermalization time of $|w_{\rm large}(n,L)\rangle$ scales with $n$, continuing to fix $n/L=10$ so as to keep both $n,L$ extensive. Operationally, we define the thermalization time as the first time when the magnitude of fluctuations in $n_{\mathtt{b}}$ drop below a fixed fraction of their initial value:
@@
Fig. 5 (c) shows the scaling of $t_{\rm th}$ with $L$, which is observed to admit an excellent fit to the predicted scaling of $\sim 2^{2n}$. As shown in Fig. 6, the distribution of thermalization times across different runs is additionally observed to be rather broad, with a standard deviation that scales approximately in the same way as $t_{\rm th}$. While this confirms that the density wave present in $|w_{\rm large}(n,L)\rangle$ relaxes on a timescale of $t_{\rm th}\sim 2^{2n}$, it does not show that all states with $\mathtt{b}$ density waves of amplitude $A$ and wavelength $q$ relax on times of order $2^{2A/q}$. This in fact cannot be true, and fast-relaxing density waves always exist. Indeed, consider the word $w_{\mathrm{small}}(n,L)$ obtained from $w_{\mathrm{large}}(n,L)$ by replacing all $\mathtt{a}$s and $\mathtt{a}^{-1}$s with $\mathtt{e}$s. The Dehn time of this word is merely ${\sf Dehn}(w_{\mathrm{small}}(n,L))\sim n^{2}$, since there are no $\mathtt{a}$s present to slow down the dynamics of the $\mathtt{b}$s (any $\mathtt{a}\mathtt{a}^{-1}$ pairs created in between the segments of the density waves have a net zero number of $\mathtt{a}$s, and are thus ineffectual at providing a slowdown). This means that the relaxation time of an initial state $|w\rangle$ carrying a $n_{\mathtt{b}}$ density wave cannot be predicted from knowledge of the conserved density alone—one must also have some knowledge about the distribution of $\mathtt{a}$’s in $|w\rangle$. We expect however that $t_{\rm th}\sim 2^{2A/q}$ for generic states containing an amplitude-$A$, wavelength-$q$ density wave. Indeed, as argued above near (77), this is simply because as long as the value of the $\mathtt{a}$-charge $\sum_{i}(|\mathtt{a}\rangle\langle\mathtt{a}|_{i}-|\mathtt{a}^{-1}\rangle\langle\mathtt{a}^{-1}|_{i})$ is nonzero in the regions between the segments of the density wave—which will almost always be true for a random density wave state in the thermodynamic limit—the $\mathtt{a}$s trapped “inside” the density wave will take an exponentially long time to escape. Note that this remark applies to a generic density wave state $|w\rangle$, regardless of whether or not $|w\rangle\in K_{e,L}$. In Fig. 7, we numerically investigate the relaxation of generic density waves by considering initial states which host density waves of momentum $q=4\pi/3L$ and amplitude $A=nq$ (with $n$ fixed at $n=L/10$), but which are otherwise random. Compared with $|w_{\rm large}(n,L)\rangle$, these random density waves (which generically lie in different $K_{g,L}$ sectors) exhibit a much broader range of thermalization times; some thermalize very quickly, while some never thermalize over our chosen simulation time window (Fig. 7 a). Nevertheless, we still observe the average thermalization timescale to scale exponentially with $n$ (Fig. 7 b). The large sample-to-sample fluctuations of $t_{\rm th}$ make it difficult to reliably extract the exact scaling behavior, but the above reasoning suggests that $t_{\rm th}$ continues to scale as $2^{2n}$ for typical initial states.
@@
Figure 7: Relaxation of $n_{\mathtt{b}}$ under stochastic circuit dynamics for initial states with random $\mathtt{b}$ density waves. The initial states are chosen to be words of the form $w_{1}\mathtt{b}^{n}w_{2}\mathtt{b}^{-n}w_{3}\mathtt{b}^{n}w_{4}\mathtt{b}^{-n}w_{5}$, where $n=L/10$ and the $w_{1,\dots,5}$ are random words containing only the characters $\{\mathtt{a},\mathtt{a}^{-1},\mathtt{e}\}$. (a) Time evolution of observable $\overline{\expectationvalue{n_{\mathtt{b}}}^{2}}$ from Eq. (89) for several independent runs with random initial states described above and the time averaging window $T=100$. The time at which the density wave collapses and the final equilibrium value of $\mkern 1.5mu\overline{\mkern-1.5mu\langle n_{\mathtt{b}}\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$ are seen to change significantly for different choices of initial state, with some runs thermalizing faster than $T$ (low-lying blue lines) and some runs not thermalizing within the whole displayed time window of $5000T$ (high-lying blue lines). (b) Thermalization times $t_{\rm th}$ of random density waves, post-selected on initial states that exhibit a drop in $\mkern 1.5mu\overline{\mkern-1.5mu\langle n_{\mathtt{b}}\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$ of at least $75\%$ during the displayed time window. $t_{\rm th}$ is determined as the median across all runs (as opposed to the mean), due to the presence of a long tail in the distribution of $t_{\rm th}$, the statistics of which are shown for $L=80$ in the inset. $t_{\rm th}$ determined in this way is seen to scale exponentially or faster with $n$.
@@ V Fragile fragmentation and the space complexity of the word problem
Our discussion in the past few sections has focused on the way in which the time complexity of the word problem enters in the thermalization times of ${\sf Dyn}_{G}$. In this section, we turn to the space complexity of the word problem. As discussed in Sec. II.2, the space complexity of the word problem is determined by the maximal amount of space required to map between two words $w,w^{\prime}\in K_{g,L}$, as diagnosed by the expansion length function ${\sf EL}_{g}(L)$ (23). When the expansion length is large, transitioning from $w$ to $w^{\prime}$ necessarily requires that $w$ first grow to be much larger than its original size before shrinking down to $w^{\prime}$. When ${\sf EL}_{g}(L)>L$, the dynamics thus lacks the spatial resources needed to connect all states which describe the same group element. In this situation, ${\sf Dyn}_{G}$ cannot act ergodically in $K_{g,L}$, and thus the $K_{g,L}$ become further fragmented. Each fragment now contains words that can be reached from some reference word $w$ by derivations that do not involve intermediate words longer than $L$. We call this phenomenon fragile fragmentation in analogy with the notion of fragile topology in band theory [40]: ${\sf Dyn}_{G}$ is said to exhibit fragile fragmentation if there are pairs of words $w,w^{\prime}$ of length $L$ such that ${\sf Dyn}_{G}$ on a system of length $L$ does not connect |$\ket{w}$ and |$\ket{w^{\prime}}$, but ${\sf Dyn}_{G}$ on a larger system of length $L^{\prime}>L$ connects the “padded” words |$\ket{w}\otimes\ket{\texttt{e}^{L^{\prime}-L}}$ and |$\ket{w^{\prime}}\otimes\ket{\texttt{e}^{L^{\prime}-L}}$. [^18] For semigroups, we would replace e with a character 00 which freely moves past other characters. A schematic illustration of this definition is given in Fig. 8. A simple example of this phenomenon that exists in higher dimensions is the jamming transition. In jammed systems, an ensemble of particles with hard core repulsive interactions can exhibit a phase transition from a low density mobile phase to a high density jammed phase. The analog of fragmentation is the limited configuration space that particles can explore in the jammed phase. When the jammed particles are given more space, their density decreases, and when it drops below a critical value, the dynamics becomes ergodic. If the extra space is subsequently removed ergodicity may again be broken, but the system may find itself in a previously inaccessible microstate. The models we discuss in this section are more drastic examples of this phenomenon: unlike the examples with hardcore particle models, the models we study exihibit jamming even in one dimension, where the analog of the critical jamming density can be polynomially or exponentially small in system size. To understand when ${\sf Dyn}_{G}$ exhibits fragile fragmentation, we need to compute the expansion length ${\sf EL}(L)$. [^19] Our notation in this discussion will assume that $G$ is a group, for which ${\sf EL}(L)\triangleq{\sf EL}_{e}(L)\gtrsim{\sf EL}_{g}(L)$ for all $g$. For general semigroup dynamics, one would need to consider the ${\sf EL}_{g}(L)$’s separately. Doing so for a general group can be rather difficult, as ${\sf EL}(L)$’s definition involves a rather complicated minimization problem. If however we already know the time complexity of the word problem—i.e. if we know the scaling of ${\sf Dehn}(L)$—it is possible to place a lower bound on ${\sf EL}(L)$ [57]. Indeed, suppose a word $w\in K_{e,L}$ has an expansion length ${\sf EL}(w)$, implying that $|w|$ is at most ${\sf EL}(w)$ during any homotopy from $w$ to $\mathtt{e}^{L}$. Then the expansion length of any derivation $D(w\rightsquigarrow\mathtt{e})$ cannot exceed the total number of length-${\sf EL}(w)$ words; if it did, there would be at least one state which appeared multiple times in $D(w\rightsquigarrow\mathtt{e})$—which implies that such a derivation cannot be of minimal length. Since the number of length ${\sf EL}(w)$ words is $(2|S|+1)^{{\sf EL}(w)}$, we thus have ${\sf Dehn}(w)\leq(2|S|+1)^{{\sf EL}(w)}$. By maximizing over all possible $w\in K_{e,L}$ and taking a logarithm, we obtain the general bound